OpenStudy (anonymous):

Which ordered pair makes the equation true? 2x – 5y = –10 (Points : 1)

OpenStudy (anonymous):

Are you given choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(5, 4) (–10, 4) (2, 5) (20, 4)

OpenStudy (anonymous):

if ur given choices just plug them in for x nd y nd c which one makes the statement true. thats the easier way

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

(5, 4) (–10, 4) (2, 5) (20, 4) The first number = x The second number = y Plug in the numbers for x and y and simplify I'll do the 3rd one for an example 2(2)-5(5)=-10 4 -25=-10 -21 = -10 Well, -21 does not equal -10, so that's not the right answer